OpenStudy (anonymous):

find the shared area between curves y=2 sin(x), y=e^(4 x), x=0, x=pi/2

myininaya (myininaya):

i would first make a rough picture of the graph of both functions together and would look for the top function and the bottom function and do \[\int\limits_{0}^{\frac{\pi}{2}}( Top-Bottom)dx\]

OpenStudy (anonymous):

thats what i did

myininaya (myininaya):

ok give me just a sec and i will do it for you and scan

myininaya (myininaya):

so it is obvious e^(4x) is above 2sin(x) lol

myininaya (myininaya):

so we have \[\frac{1}{4}e^{\frac{\pi}{2}}-\frac{1}{4}-2=\frac{1}{4}e^\frac{\pi}{2}-\frac{1}{4}-\frac{8}{4}\] \[=\frac{1}{4}e^\frac{\pi}{2}-\frac{9}{4}\]

myininaya (myininaya):

is that what you got?

myininaya (myininaya):

i;m gonna watch tv now

OpenStudy (anonymous):

thanks, but you missed one number, it not pi/2, its 2pi

myininaya (myininaya):

yea you are right it should be 1/4*e^(2pi)-9/4